### v1

@@ 5   Implementation Science
For instance, for task (A), the judge (i.e. the decision-maker), is attempting to balance public safety concerns with the well-being of the criminal defendant (i.e. decision-subject). In such a case, the preferences and priorities of the judge and the defendant are not fully aligned, and there are actions the judge can take that will be directly harmful to the defendant, but aligned with a favorable outcome for them as a judge. There are even differences in alignment between the impacts of decisions for society at large (social costs of detention or crime) vs. the judge themselves, as cases where judges release defendants who then go on to commit other crimes often receive inordinate attention and potential blowback for judges. (Kleinberg et al., 2017) refers to this as potential “omitted payoff bias” arising from this misalignment. However, for task (B), the doctor and the patient are fairly aligned stakeholders – insurance regimes such as US Medicare will not pay both the patient and healthcare provider if re-admission rate is high. Beyond insurance considerations, re-admitted patients introduce congestion costs and additional load on system, while of course patients want issues resolved without re-admission.
In addition to considering the incentives of different stakeholders relative to the impacted population, it is also worth reflecting on the institutional incentives informing the investment to introduce and integrate a prediction into the default process in the first place. What kind of cost-benefit analysis informed the process?
@@
Differences or similarities in institutional contexts across scenarios determine differences or similarities in key implementation details such as the legal ecosystem of deployment, resource availability the scale and character of the involved institutional stakeholder, the “stakes” or reversibility of a wrong decision or action, etc.
For example, the medical and financial contexts are much more regulated than, for example, criminal justice or education deployment contexts. Judicial criminal sentencing guidelines are actually much more flexible than medical clinical decision-making guidelines for diagnosis or pharmaceutical prescription since enacting discretion is part of what it means to be a judge.
@@
Because the underlying bureaucratic counterfactual policy is typically under the purview of some judicial actor, any prediction-mediated decision-making playing a similar role will often also fall under the same legal and regulatory jurisdiction.
What this means is that prediction-based decision-making is prone to similar if not more legal scrutiny than typical decision-making. In fact, the legal compatibility of evaluation and implementation is necessary for a predictor to be operable in deployment, and for the decisions made with such tools to be deemed legally valid Xiang and Raji (2019); Raghavan and Kim (2024).
@@ 6   Prescriptive Angles
Our prior analysis implies that prediction models are helpful inasmuch as they inform better decisions. In medicine, new prediction models are published daily – yet some critical appraisals question their clinical utility and point out that very few are used in clinical practice (Wynants et al., 2020; Van Smeden et al., 2022). Prediction by itself doesn’t necessarily bring about better outcomes: it’s necessarily to articulate a theory of change. This can be done via the program logic models discussed earlier in Figure 2, or by explicitly walking through how prediction information changes downstream decisions and outcomes.
@@
For some computational researchers, ADS are the closest lever in social systems to their home turf of algorithms and machine learning. But our analysis of institutional context underscores that these tools are integrated in a narrow scope of existing processes, and are thus limited in their capacity to improve them broadly. Therefore, discussions focusing on the home turf tools alone are of limited capacity to speak to broader social change. (See Abebe et al. (2020) for discussion about how computing research may nonetheless have important roles to play in social change - but usually not because of discussions about technical prediction aspects alone.)
Although debates about incrementalism and reform exist more broadly in policy (Berman and Fox, 2023), and are not just limited to ADS, discussions about ADS (and/or critiques about these discussions) can mix these together. The concern is that narrow proposals to prioritize technical improvements to the status quo might supplant or diminish (by way of technocratic authority) discussions of transformative alternatives. For example, some argue that discussions about algorithmic fairness of risk assessments are ill-scoped to advocate for criminal justice reform more broadly (Green, 2020). Another line of argument is that dedicating resources to such a technical task takes away resources from alternative reform efforts, such as policy proposals that reduce the overall rate of incarceration and/or reliance on prison systems as instruments of public safety (Hill, 2021). It’s important to be mindful of broader context and more expansive alternatives beyond prediction, even if they are not algorithmic.
Though less relevant to cases discussed earlier, it is crucial to bracket clear red lines around the appropriate deployment of social prediction to avoid inadvertent scope creep. Researchers have recently criticized efforts to predict arguably pseudoscientific physiognomic labels (e.g., “sexuality,” “criminality,” “mental abnormality”) from images, a practice that has more potential to perpetuate social harm than social benefit (Andrews et al., 2024). More subtle cases can also arise. Classification, even well-intentioned, constructs social kinds. [^23] Chicago deployed a “Strategic Subjects List” intended to predict who might be involved in violence. Later evaluations by RAND did not find benefits from the program, which by and large identified young Black men as being at-risk for violence. The original researchers who created the algorithm emphasize that the initial interest was to provide social service outreach. But by and large, SSL was involved more in arrests than supportive social services, and introduced stigmatic harms. Potential stigmatic harms ought to be weighed against any potential benefits.
Prediction is one of many paths that can improving outcome where technical researchers can contribute expertise. In practice, there exist other alternatives for broader consideration. To give an example of alternative decision processes,we discuss lotteries (including partial, weighted, or tiered lotteries), where interventions are randomly assigned among a possibly pre-screened pool of eligible individuals. This approach acknowledges both the limits of prediction and the normative concern that highly targeted interventions can unintentionally reinforce existing inequalities by systematically concentrating harms or benefits in the targeted subpopulation. In contrast, lotteries could provide a fairer (Stone, 2007) and more transparent mechanism for resource allocation, as demonstrated in healthcare (White and Angus, 2020), housing (Arnosti and Shi, 2020), and research funding (Horbach et al., 2022). Though some work studies when lotteries are aligned with social preference weightings (Heidari et al., 2018), further theoretical and empirical work is needed to better understand the properties, feasibility, and potential impacts of lotteries in comparison with prediction-based targeting methods. Analogously, technical research can quantify and illuminate operating characteristics of other alternative decision processes beyond prediction.
@@
In the end, it is important to critically evaluate whether algorithmic development is necessary and the best way to advance the social good in context; sometimes the best solution is not an algorithmic one.
@@
Changing the estimand:
When algorithmic solutions are desirable, the appropriate tools may not always be prediction models. Section 3 outlines these alternatives in detail, including their technical formulations. Changing the problem formulation, changing the estimand, or introducing additional desiderata can all be crucial tools for better aligning ADS with outcomes of interest. Identifying causal determinants (a.k.a. “causal predictors”) of predicted outcomes can highlight targets for intervention. Optimizing individualized treatment rules give utility-optimal decisions in the face of treatment effect heterogeneity. Estimating counterfactual risk models can better estimate predictive risk in view of static interventions (Keogh and Van Geloven, 2024; Coston et al., 2020).
@@
Revisiting data collection
Researchers working on ADS systems often treat the data available as fixed, relying on existing types of administrative records to train predictive models. The design of these predictive models is therefore a consequence of the type of data available for training. This is generally not first-best, where researchers first prospectively decide what comprises useful data, and then collect that data to create a predictive model. Yet there may be many cases where researchers should consider whether it is feasible to collect or access additional data, and what the costs and benefits of such a process would involve. We suggest that explicitly considering these questions of what data would improve a model, whether it exists, and what it would take to collect it, is critical for improving predictive models.
When the true outcome of interest is not collected in existing data, it is worth evaluating whether it is feasible to do so, what are the potential benefits, and costs. Indeed, the social cost of tracking individual trajectories can be extreme. Trading off these costs might suggest triangulation based on population statistics.
@@
While we have focused on welfarist accounts of the utility of algorithms, understanding their deployment and procedural transparency/opacity is crucial. Algorithms are often painted in context as either opaque or transparent, depending on context and use. Some argue that algorithms are ‘black boxes,’ machines whose internal logic is impossible to parse or explain (Pasquale, 2015). By other accounts, algorithms offer a unique opportunity to move away from subjective and often biased human judgments and towards rule-based systems that are easier to vet and understand (Mullainathan, 2019). In many consequential domains, procedural due process protections are required. This can take the form of appeals processes or human review, which can be important for contesting incorrect algorithmic inputs.
The affordances and requirements of algorithms change significantly depending on context. An algorithm can simultaneously make some aspects of a decision more transparent (e.g., the ordering of inferred risks across individuals) while making others more opaque (e.g., the process through which risk scores are inferred). The first step towards making normative conclusions about whether algorithmic decisions improve outcomes in a social systems is to measure their full effects within a specific context or use. Algorithmic systems need to be vetted for both efficacy, safety, and procedural effects, especially in light of complex processes already in place for performing similar functions (Laufer and Nissenbaum, 2023).
@@
Developing mechanisms and governance structures for performance management is necessary when it is not already present. Even with existing frameworks, it’s impossible to anticipate all harms ahead of time. Experimenting with new structures for continual performance monitoring is crucial. Although incident report databases such as those used by vaccines (U.S. Department of Health and Human Services, ) have some issues with implementation (e.g., selective reporting, lack of knowledge by users), they can provide a feedback mechanism for deployers.
@@ Appendix A Additional discussion on Background
Example 9 (Dropout Prediction).
 [^27] Adapted in part from Wisconsin Department of Public Instruction (2025) and Perdomo et al. (2023). [Wisconsin Public Schools Dropout Prediction]
Dropout Early Warning System (DEWS) (Knowles, 2015) was a risk scoring system designed and implemented by the Wisconsin Department of Public Instruction (DPI) from 2012 to 2023 (Wisconsin Department of Public Instruction, 2025). The system provides educators with risk scores for all middle school students (grades 6-9), generating these predictions using over 40 student features. Features include demographic and socioeconomic information (e.g., race, gender, family income), academic performance (e.g., scores on state standardized exams), as well as community-level statistics (e.g., percent of cohort that is non-White and school size). The system has two main outputs: 1) the DEWS risk category (low, moderate or high) and 2) the DEWS score (an estimated probability of on-time graduation that takes continuous values between 0 and 1). Models are fit by empirical risk minimization on a dataset of a historical cohort of students, and are updated every year. While DPI maintains that DEWS dropout predictions “can lead to critical interventions that prevent students from actually dropping out.”, there has been no prescriptive guidelines provided to schools on how to act on the DEWS scores (Feathers, 2023) and no systematic data collection on the actual educational interventions applied in response to DEWS.
@@ Appendix B Additional discussion on Model Design
Local differences in capacity lead to a “tie-breaking” design where local offices provided mandatory services for highest-risk claimants, randomizing at the discrete score value that satisfies the capacity constraint. The use of mandatory services was thought to improve moral hazard. While the treatment group overall returned earlier to work, improving total earnings, the timing of these early returns coincided with the onset of mandatory work requirements. The variability in local offices allowed for assessing treatment effects in quantiles of risk, finding evidence against a “common effect” assumption that justifies prioritization by expected duration (in the absence of the program). That is, $E[Y(1)-Y(0)|Y(0)]$, as assessed via low, moderate, and high scores, was highly nonlinear in the predicted risk, with the greatest benefits accruing to those with moderate risk of exhausting benefits, overall providing “evidence that allocating the treatment on the basis of the expected duration of UI benefit receipt may not represent an optimal strategy.”
@@
So far, we have assumed that the decision maker knows a specific loss $\ell$ that they wish to minimize. A recent line of work in machine learning seeks predictors which are near-optimal for decision making with respect to some class of potential downstream losses.282828 Technically, these formulations tend to be closely related to multicalibration, a solution concept which requires predictors to be conditionally calibrated for each of a (potentially large) number of subgroups. One example is the notion of an omnipredictor. Roughly, a omnipredictor relative to a set of loss functions $\mathcal{L}$ and a comparison class of models $\mathcal{C}$ guarantees that for every loss in $\mathcal{L}$, its predictions suffice to make decisions nearly as good as any model in $\mathcal{C}$. A related notion is decision calibration, introduced by (Zhao et al., 2021). Decision calibration requires that for any loss $\ell$ in a set $\mathcal{L}$ and any alternative $\ell^{\prime}\in\mathcal{L}$ the decision maker obtains lower loss (in expectation) by selecting $D\in\arg\min\mathbb{E}_{Y\sim\hat{p}(Y|X)}[\ell(D,Y)]$ than $D^{\prime}\in\arg\min\mathbb{E}_{Y\sim\hat{p}(Y|X)}[\ell^{\prime}(D,Y)]$. Intuitively, this can be understood as requiring that the decision maker cannot gain by “post-processing” the predictions; they are incentivized to best-respond as if the predictions $\hat{p}$ are in fact correct. Similar ideas were recently generalized by (Noarov et al., 2023) beyond the scalar case to predicting a multidimensional state followed by optimization.
@@
While interference appears the most “obviously” vulnerable assumption for the social world, other assumptions are less obviously vulnerable, such as modularity of interventions, “well-defined interventions”, and consistency. Some argue that “well-defined interventions” are inherently conservative, in the sense of limiting interventions to local perturbations around the status quo (Stevenson, 2023).323232 Stevenson (2023) contends that in general, the kinds of interventions that lend themselves to high-quality program evaluations and pilots reflect the constraints of programmatic activities, rather than pursuing structural social change which might be more transformative in improving outcomes. More specifically, (Stevenson, 2023) considers the mixed evidence base for interventions in the criminal-legal system and finds that overall, treatment effects of published studies are usually not statistically significant, and that studies that show promise in one location struggle to scale to other locations. On the other hand, structural social change can be inherently ill-suited to the framework of causal comparison, which seeks to isolate differences in context and setting to that of the intervention alone, “holding fixed everything but the intervention”.

### v2

@@ 5  Implementation Science
For instance, for task (A), the judge (i.e. the decision-maker), is attempting to balance public safety concerns with the well-being of the criminal defendant (i.e. decision-subject). In such a case, the preferences and priorities of the judge and the defendant are not fully aligned, and there are actions the judge can take that will be directly harmful to the defendant, but aligned with a favorable outcome for them as a judge. There are even differences in alignment between the impacts of decisions for society at large (social costs of detention or crime) vs. the judge themselves, as cases where judges release defendants who then go on to commit other crimes often receive inordinate attention and potential blowback for judges. (Kleinberg et al., 2017) refers to this as potential “omitted payoff bias” arising from this misalignment. However, for task (B), the doctor and the patient are fairly aligned stakeholders – insurance regimes such as US Medicare will not pay both the patient and healthcare provider if re-admission rate is high. Beyond insurance considerations, re-admitted patients introduce congestion costs and additional load on system, while of course patients want issues resolved without re-admission. Not only do individual decision-makers face these incentives, but also the institutions that decided to introduce ADS in the first place.
@@
Differences or similarities in institutional contexts across scenarios give rise to key contrasts or similarities in implementation design such as the legal ecosystem, the scale of the decision-making, the stakes, recourse or reversibility of a wrong decision or action, among others. For example, the medical and financial contexts are much more regulated than, for example, criminal justice or education deployment contexts. Judicial criminal sentencing guidelines have substantial room for inherent discretion, compared to medical clinical decision-making guidelines for diagnosis or prescription.
@@
Prediction-mediated decision-making falls under the same legal and regulatory jurisdiction as the deploying institution, and is often prone to similar if not more legal scrutiny than typical decision-making. In fact, the legal compatibility of evaluation and implementation is necessary for a predictor to be operable in deployment and legally valid Xiang and Raji (2019); Raghavan and Kim (2024).
@@ 6  Prescriptive Angles
Our prior analysis implies that prediction models are helpful inasmuch as they inform better decisions. In medicine, new prediction models are published daily – yet some critical appraisals question their clinical utility and point out that very few are used in clinical practice (Wynants et al., 2020; Van Smeden et al., 2022). Prediction by itself doesn’t necessarily bring about better outcomes: it’s necessarily to articulate mechanisms for how they would do so, walking through how prediction information should change downstream decisions and outcomes..
@@
Overall, we have focused on socio-technical questions regarding evaluating ADS. But our analysis of institutional context underscores that these tools are integrated in a narrow scope of existing processes, and are thus limited in their capacity to improve them broadly. Technical concerns re: developing, evaluating, and improving ADS are necessarily just a small part of broader discussions of social change and improvement. (See Abebe et al. (2020) for discussion about how computing research may nonetheless have important roles to play in social change - but rarely from discussing technical prediction aspects alone.)
It’s important to be consider expansive alternatives beyond prediction: advocating for more resources, improving intervention effectiveness, and many others. One example alternative is that of lotteries (including partial, weighted, or tiered lotteries), where interventions are randomly assigned among a possibly pre-screened pool of eligible individuals, pose an alternative approach. Lotteries could provide a fairer (Stone, 2007) and more transparent mechanism for resource allocation, as demonstrated in healthcare (White and Angus, 2020), housing (Arnosti and Shi, 2020), and research funding (Horbach et al., 2022).
@@
In the end, it is important to critically evaluate whether algorithmic development is necessary; sometimes the best solution is not an algorithmic one.
@@
Changing the estimand: When algorithmic solutions are desirable, the appropriate tools may not always be prediction models, but interventional or decision-aware approaches. Section 3 outlines these alternatives in detail, including their technical formulations. Changing the problem formulation, changing the estimand, or introducing additional desiderata can all be crucial tools for better aligning ADS with outcomes of interest. Alternative approaches include identifying causal determinants as candidates for intervention, optimizing treatment rules, or estimating counterfactual risk models.
@@
Revisiting data collection Researchers working on ADS systems often treat the data available as fixed, relying on existing types of administrative records to train predictive models. Yet there may be many cases where researchers should consider whether it is feasible to collect or access additional data, and what the costs and benefits of such a process would involve. When the true outcome of interest is not collected in existing data, it is worth evaluating whether it is feasible to do so, what are the potential benefits, and costs. Indeed, the social cost of tracking individual trajectories can be extreme. Trading off these costs might suggest triangulation based on population statistics.
@@
While we have focused on welfarist accounts of the utility of algorithms, understanding their deployment and procedural transparency/opacity is crucial. Algorithms are often painted in context as either opaque black boxes (Pasquale, 2015) or more transparent alternatives to the black-box of human decisions (Mullainathan, 2019), depending on context and use . In many consequential domains, procedural due process protections are required. This can take the form of appeals processes or human review, which can be important for contesting incorrect algorithmic inputs.
The affordances and requirements of algorithms change significantly depending on context. An algorithm can simultaneously make some aspects of a decision more transparent (e.g., the ordering of inferred risks across individuals) while making others more opaque (e.g., the process through which risk scores are inferred). Algorithmic systems need to be vetted for both efficacy, safety, and procedural effects, especially in light of complex processes already in place for performing similar functions (Laufer and Nissenbaum, 2023).
@@
Developing mechanisms and governance structures for performance management is necessary when it is not already present. It’s still impossible to anticipate all harms ahead of time, so experimenting with new structures for continual performance monitoring is crucial. Although incident report databases such as those used by vaccines (U.S. Department of Health and Human Services, ) have some issues with implementation (e.g., selective reporting, lack of knowledge by users), they can provide a feedback mechanism for deployers.
@@ Appendix A Additional discussion on Background
Example 8 (Wisconsin Public Schools Dropout Prediction).
Dropout Early Warning System (DEWS) (Knowles, 2015) was a risk scoring system designed and implemented by the Wisconsin Department of Public Instruction (DPI) from 2012 to 2023 (Wisconsin Department of Public Instruction, 2025). The system provides educators with risk scores for all middle school students (grades 6-9), generating these predictions using over 40 student features. Features include demographic and socioeconomic information (e.g., race, gender, family income), academic performance (e.g., scores on state standardized exams), as well as community-level statistics (e.g., percent of cohort that is non-White and school size). The system has two main outputs: 1) the DEWS risk category (low, moderate or high) and 2) the DEWS score (an estimated probability of on-time graduation that takes continuous values between 0 and 1). Models are fit by empirical risk minimization on a dataset of a historical cohort of students, and are updated every year. While DPI maintains that DEWS dropout predictions “can lead to critical interventions that prevent students from actually dropping out.”, guidelines provided to schools on how to act on the DEWS scores leave out key model information and provide insufficient guidance for interpreting these scores (Feathers, 2023). There has been no systematic data collection on the actual educational interventions applied in response to DEWS.
@@ Appendix B Additional discussion on Model Design
Local differences in capacity lead to a “tie-breaking” design where local offices provided mandatory services for highest-risk claimants, randomizing at the discrete score value that satisfies the capacity constraint. The use of mandatory services was thought to improve moral hazard. While the treatment group overall returned earlier to work, improving total earnings, the timing of these early returns coincided with the onset of mandatory work requirements. The variability in local offices allowed for assessing treatment effects in quantiles of risk, finding evidence against a “common effect” assumption that justifies prioritization by expected duration (in the absence of the program). That is, E[Y(1)Y(0)|Y(0)]E[Y(1)-Y(0)|Y(0)], as assessed via low, moderate, and high scores, was highly nonlinear in the predicted risk, with the greatest benefits accruing to those with moderate risk of exhausting benefits, overall providing “evidence that allocating the treatment on the basis of the expected duration of UI benefit receipt may not represent an optimal strategy.”
@@
So far, we have assumed that the decision maker knows a specific loss $\ell$ that they wish to minimize. A recent line of work in machine learning seeks predictors which are near-optimal for decision making with respect to some class of potential downstream losses. [^32] Technically, these formulations tend to be closely related to multicalibration, a solution concept which requires predictors to be conditionally calibrated for each of a (potentially large) number of subgroups. One example is the notion of an omnipredictor. Roughly, a omnipredictor relative to a set of loss functions $\mathcal{L}$ and a comparison class of models $\mathcal{C}$ guarantees that for every loss in $\mathcal{L}$, its predictions suffice to make decisions nearly as good as any model in $\mathcal{C}$. A related notion is decision calibration, introduced by (Zhao et al., 2021). Decision calibration requires that for any loss $\ell$ in a set $\mathcal{L}$ and any alternative $\ell^{\prime}\in\mathcal{L}$ the decision maker obtains lower loss (in expectation) by selecting $D\in\arg\min\mathbb{E}_{Y\sim\hat{p}(Y|X)}[\ell(D,Y)]$ than $D^{\prime}\in\arg\min\mathbb{E}_{Y\sim\hat{p}(Y|X)}[\ell^{\prime}(D,Y)]$. Intuitively, this can be understood as requiring that the decision maker cannot gain by “post-processing” the predictions; they are incentivized to best-respond as if the predictions $\hat{p}$ are in fact correct. Similar ideas were recently generalized by (Noarov et al., 2023) beyond the scalar case to predicting a multidimensional state followed by optimization.
@@
In this vein, performative prediction (Perdomo et al., 2020) is a recent line of work (Mendler-Dünner et al., 2022; Miller et al., 2021; Hardt et al., 2022; Kim and Perdomo, 2023; Perdomo, 2025) that explicitly aims to analyze the feedback loops between predictive algorithms and the social world. It is motivated by the fact that whenever we make predictions about social events, these predictions shape people’s expectations and actions in ways that end up influencing the likelihood of the the predicted outcome and the distribution of observed features. For instance, predicting that someone is at risk of a heart attack influences their lifestyle choices in ways that shape their future outcomes. Formally, performative prediction is a learning theoretic framework that, amongst other things, has been used to understand different kinds of social prediction dynamics such as repeated retraining. See Hardt and Mendler-Dunner (2025) for an overview of work in this space.
@@
While interference appears the most “obviously” vulnerable assumption for the social world, other assumptions are less obviously vulnerable, such as modularity of interventions, “well-defined interventions”, and consistency. Some argue that “well-defined interventions” are inherently conservative, in the sense of limiting interventions to local perturbations around the status quo (Stevenson, 2023). [^36] Stevenson (2023) contends that in general, the kinds of interventions that lend themselves to high-quality program evaluations and pilots reflect the constraints of programmatic activities, rather than pursuing structural social change which might be more transformative in improving outcomes. More specifically, (Stevenson, 2023) considers the mixed evidence base for interventions in the criminal-legal system and finds that overall, treatment effects of published studies are usually not statistically significant, and that studies that show promise in one location struggle to scale to other locations. On the other hand, structural social change can be inherently ill-suited to the framework of causal comparison, which seeks to isolate differences in context and setting to that of the intervention alone, “holding fixed everything but the intervention”.
@@ Appendix C Additional discussion on Evaluation Science
Challenges and Open Problems for RCTs of ADS in Social Settings